OpenStudy (anonymous):

Los Angeles is about 385 miles from San Francisco. How far apart would the cities be on a map with a scale of 1 in. = 15 mi?

OpenStudy (mathmale):

I'd treat 1 inch / 15 miles as a conversion factor: \[(385 miles)*(1 inch / 15 miles) = ??\]

OpenStudy (anonymous):

385 miles * 1 inch devided by 15 miles>

OpenStudy (anonymous):

?*

OpenStudy (anonymous):

that would be 25.6

OpenStudy (mathmale):

\[\frac{ (385 miles)*(1 inch) }{ 15 miles }=?\]

OpenStudy (mathmale):

25,7 inches, approximately.

OpenStudy (mathmale):

So, 385 miles (actual distance) is represented by 25.7 inches on this map.

OpenStudy (anonymous):

Thanks, I didnt understand how to solve it

OpenStudy (mathmale):

U OK now?

OpenStudy (anonymous):

yeah